# A systematic study of AGN feedback in a disk galaxy II: MACER prediction of X-ray surface brightness profile and comparison with eROSITA observations

### 23 Mar 2026 v1

@@ IV.3 The importance of AGN Feedback
To investigate the role of AGN feedback in affecting the X-ray surface brightness profile, we computed the prediction of the noAGN model from Paper I. Figure 4 compares the predictions of the noAGN model (dashed curves) with those of the Fiducial model (solid curves) against two observational datasets. In the left panel, when compared with the stacked profile of Milky Way-mass galaxies from Z24, the noAGN model underpredicts the X-ray surface brightness by nearly an order of magnitude from a few kiloparsecs out to $\sim 100\,\mathrm{kpc}$. In contrast, the right panel shows that, relative to the stacked measurements of He and Li (2026), the noAGN model is lower than the Fiducial prediction by a factor of $\sim 2$–3 only around $R\sim 30\,\mathrm{kpc}$, while at larger radii ($R\gtrsim 30\,\mathrm{kpc}$) the two models yield comparable surface brightness profiles.
@@
To understand the physical reason, we show the time-averaged radial profiles of the number density and temperature of the CGM gas for the Fiducial run and the noAGN run in the left and middle panels of Fig. 5. The AGN feedback suppresses the central gas density by a factor of a few within $\sim 100\,\mathrm{kpc}$, but increases the temperature by a factor of a few, from below $10^{6}\,\mathrm{K}$ to $10^{6}$$10^{7}\,\mathrm{K}$ throughout the galaxy. The time-averaged radial entropy profile is correspondingly enhanced: with AGN feedback active, the entropy is higher by nearly two orders of magnitude at essentially all radii compared to runs where AGN feedback is switched off. The results clearly indicate that the physical properties of the CGM are regulated by AGN feedback: in the absence of feedback, the model predictions lie systematically below the observational data (left panel of Fig. 4), demonstrating that AGN feedback plays a crucial role.
Note that at $R\simeq 100\,\mathrm{kpc}$, the noAGN run has a higher gas density, and the emissivity at $T\simeq 2\times 10^{6}\,\mathrm{K}$ exceeds that at $10^{7}\,\mathrm{K}$, yet its surface brightness is lower than that in the Fiducial model. This is a projection effect: the X-ray surface brightness scales with the line-of-sight emission measure:
$S_{X}\propto\int n_{e}^{2}\,\Lambda(T,Z)\,dl.$
Physically, gas near $2\times 10^{6}\,\mathrm{K}$ harbors abundant partially ionized metals (e.g., Fe, O, Si) that yield strong line emission via numerous electronic transitions, whereas at $10^{7}\,\mathrm{K}$, these elements are nearly fully stripped, suppressing line emission and leaving the much less efficient free–free continuum as the dominant process.
@@
The fact that the noAGN model deviates from the Fiducial model in the right panel of Fig. 4 only within a narrow radial range around $R\sim 30\,\mathrm{kpc}$, yet yields comparable X-ray surface brightness at larger radii, can be understood as the combined consequence of gas density, temperature, and metallicity on the X-ray emissivity. In the noAGN run, the gas density at these radii is higher while the temperature is lower than in the Fiducial run. A lower temperature tends to reduce the bremsstrahlung (free-free) continuum, but once the gas cools to $T\sim 10^{6}\,\mathrm{K}$, the cooling function $\Lambda(T,Z)$ enters a regime dominated by metal-line emission and becomes highly sensitive to the assumed metallicity. This difference in the relative importance of line versus continuum emission is also why the left and right panels of Fig. 4 behave differently. In the comparison with He and Li (2026) (right panel), a solar metallicity ($Z=Z_{\odot}$) is adopted, which substantially enhances metal-line emissivity around $T\sim 10^{6}\,\mathrm{K}$. As a result, the increase in line emission (together with the higher density) compensates for the reduced continuum at lower temperatures, so the predicted surface brightness of the noAGN model at $R\gtrsim 30$$40\,\mathrm{kpc}$ is boosted and becomes comparable to that of the Fiducial model. By contrast, in the comparison where a sub-solar metallicity is assumed (e.g., $Z=0.3\,Z_{\odot}$; left panel), the line-emission boost is much weaker, so the temperature reduction leads to a systematically lower emissivity and hence a markedly fainter surface brightness in the noAGN model over a broader radial range.
V Summary
@@
Most recently, we have studied the evolution of a disk galaxy using MACER, a framework designed to study the effects of AGN feedback on the evolution of a single galaxy. The key features of MACER are that the Bondi radius is resolved so that the black hole accretion rate is reliably determined, and that state-of-the-art AGN physics has been incorporated into the model. The simulation data have been analyzed in detail, and several scientific problems have been addressed and compared with observations, including the AGN duty cycle, the correlation between black hole accretion rate and star formation rate, the galaxy quenching mechanism, and so on. In all cases, good consistency with observations has been obtained, and the results will be presented in a series of papers.
@@
While Paper I of this series focuses on the global overview of the simulation, as the second paper of this series, we compute the surface brightness profile predicted by the Fiducial model presented in Paper I and compare it with the recent eROSITA measurements of the X-ray surface brightness of the circumgalactic medium (CGM) around Milky Way-mass galaxies. All model parameters are identical to those adopted in Paper I and no additional free parameters are introduced. We compare our model predictions with two independent eROSITA-based observational studies: Z24 and He and Li (2026). Although both studies probe the X-ray emission from the hot CGM, they differ in galaxy sample selection, distance (or redshift) range, and analysis methodology. Z24 measures the mean CGM X-ray surface brightness profile using a large stacked sample of Milky Way–mass galaxies over a broad redshift range, whereas He and Li (2026) focuses on a sample of nearby $\rm L$$^{*}$ galaxies, enabling a complementary view of the CGM at smaller physical scales. Our main results are summarized as follows.
@@
Since there is no metallicity information in the Paper I simulations, we assume four representative metallicities of $Z=0.02\,Z_{\odot}$, $Z=0.3\,Z_{\odot}$, $Z=0.5\,Z_{\odot}$, and $Z=1.0\,Z_{\odot}$. We find that the mean simulated surface-brightness profiles are in good agreement with the stacked measurements of Z24 over a wide range of radii, out to $\sim 100\,\mathrm{kpc}$. The model predictions are also consistent with the results of He and Li (2026) at radii $r\gtrsim 20\,\mathrm{kpc}$, extending to $r\sim 120\,\mathrm{kpc}$ (Figure 1). Overall, this level of agreement indicates that the eROSITA-detected X-ray emission is predominantly produced by thermal radiation from hot gas, with no requirement for a significant nonthermal contribution, in line with the conclusions of He and Li (2026). In particular, He and Li (2026) shows that the stacked 10–50 kpc spectrum is well described by a thermal plasma plus a power-law component, where the power-law is consistent with residual emission from discrete sources (unresolved stellar populations and PSF-scattered photons) and leaves little room for an additional inverse-Compton/CR-related continuum; hence, a CR-dominated interpretation of the diffuse halo emission is strongly disfavored. Given the limited spectral quality of the current data, however, a small nonthermal (CR-related) contribution cannot be strictly excluded. Deeper eROSITA data from later all-sky surveys (e.g., eRASS4) will improve the spectral constraints and may enable measurements to larger radii.
@@
Y.Z., F.Y., and S.J. are supported by the NSF of China (grants 12192220, 12192223, 12522301, 12133008, and 12361161601), the China Manned Space Program (grants CMS-CSST-2025-A08 and CMS-CSST-2025-A10), the National Key R&D Program of China No. 2023YFB3002502, and the National SKA Program of China (No. 2025SKA0130100). T. F. is supported by the National SKA Program of China No. 2025SKA0150103 and the National Natural Science Foundation of China under Nos. 12550002, 12133008, 12221003, and 11890692. We acknowledge the science research grants from the China Manned Space Project with grant Nos. CMS-CSST-2021-A04 and CMS-CSST-2025-A10. Numerical calculations were run on the CFFF platform of Fudan University, the supercomputing system at the Supercomputing Center of Wuhan University, and the High Performance Computing Resource at the Core Facility for Advanced Research Computing at Shanghai Astronomical Observatory.

### 01 May 2026 v2

@@ IV.3 The importance of AGN Feedback
To investigate the role of AGN feedback in affecting the X-ray surface brightness profile, we computed the prediction of the noAGN model from Paper I. Figure 4 shows the comparison between the noAGN model (dashed curves) and the Fiducial model (solid curves) against two observational datasets. In the left panel, when compared with the stacked profile of Milky Way-mass galaxies from Z24, the noAGN model underpredicts the X-ray surface brightness by nearly an order of magnitude from a few kiloparsecs out to $\sim 100\,\mathrm{kpc}$. In contrast, the right panel shows that, over $R\sim 20$$30\,\mathrm{kpc}$, the noAGN model falls below the Fiducial prediction only by a factor of $\sim 2$–3 whereas at larger radii ($R\gtrsim 30\,\mathrm{kpc}$) the predictions of the two models are comparable [^2] Given the observational limitations discussed in §IV.1, our comparison mainly focuses on the region of $R\gtrsim 20\,\mathrm{kpc}$.. The results clearly indicate that the physical properties of the CGM are regulated by AGN feedback: in the absence of feedback, the model predictions lie systematically below the observational data, especially in the case of the left panel of Fig. 4, demonstrating that AGN feedback plays a crucial role.
@@
To understand the physical reason of the results shown in Fig. 4, we show the time-averaged radial profiles of the number density and temperature of the CGM gas for the Fiducial run and the noAGN run in the left and middle panels of Fig. 5. The AGN feedback suppresses the gas density by a factor of a few to more than an order of magnitude within $\sim 100\,\mathrm{kpc}$, but increases the temperature by a factor of a few, from below $10^{6}\,\mathrm{K}$ to $10^{6}$$10^{7}\,\mathrm{K}$ throughout the galaxy. The time-averaged radial entropy profile is correspondingly enhanced; with AGN feedback active, the entropy is higher by about one order of magnitude at essentially all radii compared to runs where AGN feedback is switched off.
The X-ray surface brightness profile presented in Fig. 4 can be interpreted as follows. The surface brightness $S_{X}$ is proportional to the emission measure integrated along the line of sight:
$S_{X}\propto\int n_{e}^{2}\,\Lambda(T,Z)\,dl,$
where $\Lambda(T,Z)$ includes both bremsstrahlung continuum and metal-line emission, the latter is sensitive to metallicity. Physically, gas at $T\sim 2\times 10^{6}\,\mathrm{K}$ contains abundant partially ionized metals (e.g., Fe, O, and Si), which produce strong line emission through numerous electronic transitions, whereas at $T\sim 10^{7}\,\mathrm{K}$ these elements are nearly fully stripped, so the line contribution is strongly suppressed and the free–free continuum becomes dominant. In the left panel of Fig. 4, the Fiducial run predicts noticeably higher surface brightness than the noAGN run, even though the latter yields a higher gas density. This apparent discrepancy arises from three combined effects. The first one is that the higher temperature in the Fiducial run enhances the bremsstrahlung emission. The second reason is that the subsolar metallicity ($Z=0.3\,Z_{\odot}$) assumed in this panel reduces the relative contribution of line emission, making the temperature-sensitive bremsstrahlung more dominant. The third is the projection effect. The observed brightness at a given projected radius integrates emission along the entire line of sight. Significant contributions come from larger physical radii, where the densities in the two runs become similar, but the temperature (and thus the emissivity) in the Fiducial run remains substantially higher.
@@
In the right panel, however, the difference in predicted surface brightness between the two runs is less pronounced compared to the left panel. The temperature in the noAGN run $T\sim 10^{6}\,\mathrm{K}$, so the cooling function $\Lambda(T,Z)$ enters a regime dominated by metal-line emission and becomes highly sensitive to the assumed metallicity. In the comparison with He and Li (2026) (right panel), a solar metallicity ($Z=Z_{\odot}$) is adopted, which substantially enhances metal-line emissivity around $T\sim 10^{6}\,\mathrm{K}$. As a result, the increase in line emission (together with the higher density) compensates for the reduced continuum at lower temperatures, so the predicted surface brightness of the noAGN model is boosted and becomes comparable to that of the Fiducial model.
We have also examined whether a predominantly thermal CGM in our model would necessarily predict an excessively strong tSZ signal. For the Fiducial model, the time-averaged integrated tSZ signal (with the 7.8–8.2 Gyr high-AGN-luminosity phase excluded) is $Y_{200}^{\mathrm{sph}}\sim 2.4\times 10^{-8}\ {\rm arcmin}^{2}$ when the galaxy is placed at a distance of 500 Mpc, with temporal scatter spanning a factor of $\sim 4$ across different epochs. This stays slightly below the current observational upper limits of tSZ stacking by Das et al. (2025), while the mean value of $Y_{500}^{\mathrm{sph}}$ predicted by our simulations is $\sim 7.4\times 10^{-8}\ {\rm arcmin}^{2}$, systematically lower than the observational value of $Y_{500}^{\mathrm{sph}}\sim 10^{-7}\,\mathrm{arcmin}^{2}$ reported by Bregman et al. (2022). During peak AGN outburst phases, $Y_{200}^{\mathrm{sph}}$ can reach $\sim 10^{-6}\,\mathrm{arcmin}^{2}$; such peak outburst values can exceed observational medians, but the duty-cycle-averaged signal remains broadly consistent within uncertainties. This result indicates that a thermal halo does not necessarily imply an overly large tSZ signal, and provides a complementary perspective to the non-thermal interpretation (Ponnada et al., 2026).
V Summary and discussion
@@
Most recently, we have studied the evolution of a disk galaxy using MACER, a framework designed to study the effects of AGN feedback on the evolution of a single galaxy. The simulation data have been analyzed in detail, and several scientific problems have been addressed and compared with observations, including the AGN duty cycle, the correlation between black hole accretion rate and star formation rate, the galaxy quenching mechanism, and so on. The results will be presented in a series of papers.
@@
While Paper I in this series presents a global overview of the simulation, this second paper focuses on the surface brightness profile predicted by the Fiducial model introduced in that work. Specifically, we compute the X-ray surface brightness profile and compare it with recent eROSITA measurements of the circumgalactic medium (CGM) surrounding Milky Way–mass galaxies. All model parameters are identical to those adopted in Paper I, and no additional free parameters are introduced. We compare our model predictions with two independent observational studies based on eROSITA data: Z24 and He and Li (2026). Although both studies probe X-ray emission from the hot CGM, they differ in galaxy sample selection, redshift range, and analysis methodology. Z24 derives the mean CGM X-ray surface brightness profile from a large stacked sample of Milky Way–mass galaxies spanning a broad redshift range, whereas He and Li (2026) focuses on a sample of nearby $\rm L$$^{*}$ galaxies, providing a complementary view of the CGM on smaller physical scales. Our main results are summarized as follows.
@@
Since there is no metallicity information in the Paper I simulations, we assume four representative metallicities of $Z=0.02\,Z_{\odot}$, $Z=0.3\,Z_{\odot}$, $Z=0.5\,Z_{\odot}$, and $Z=1.0\,Z_{\odot}$. We find that the mean simulated surface-brightness profiles are in good agreement with the stacked measurements of Z24 over a wide range of radii, out to $\sim 100\,\mathrm{kpc}$. The model predictions are also consistent with the results of He and Li (2026) at radii $r\gtrsim 20\,\mathrm{kpc}$, extending to $r\sim 120\,\mathrm{kpc}$ (Figure 1). This agreement indicates that the eROSITA-detected X-ray emission could be predominantly produced by thermal radiation from hot gas, with no requirement for a significant nonthermal contribution, in line with the conclusions of He and Li (2026). We caution, however, that this consistency does not by itself uniquely establish a thermal origin: uncertainties in the assumed initial CGM gas mass, together with the absence of cosmic rays in the current simulations, leave open the possibility that a nonthermal component contributes at some level.
@@
We would like to point out two caveats of our work, which should be kept in mind when interpreting the comparison. First, since the simulations in Zou et al. (2026) are performed for a single galaxy, in the present work, we approximate a galaxy ensemble with a time ensemble; that is, we stack outputs from different time steps of the same galaxy to mimic stacking different galaxies. Although real galaxies vary in dark matter halo properties, bulge-to-disk ratios, and other structural parameters, we contend that this approximation does not pose a significant problem for the present study. The X-ray surface brightness profile depends directly on gas density and temperature, the effects of other galaxy properties enter mainly through their influence on these two quantities. Given the limited diversity in structural properties across the observed sample, we expect galaxy mass and, in particular, AGN feedback to play a more important role in determining the X-ray surface brightness profile. For example, at different epochs, AGN feedback can change the gas content of a galaxy by as much as three orders of magnitude, which would translate into an approximately six-order-of magnitude variation in the X-ray surface brightness. In this sense, our treatment should capture the zeroth-order behavior of the observational stacks.
Second, the Zou et al. (2026) simulations span $\sim$ 12 Gyr, whereas the observations focus on low-redshift galaxies. We note that the Zou et al. (2026) simulations are still somewhat idealized. As we emphasize in that paper, they do not account for important physical processes such as galaxy mergers and the evolution of the galactic gravitational potential driven by the cosmological growth of dark matter halos, thus the $\sim$12 Gyr evolution should not be directly compared with observations spanning different redshifts. Our goal is to isolate and investigate the key physical mechanisms that regulate galaxy evolution, rather than to obtain a precise cosmological evolution of the galaxies. Importantly, the initial gas density adopted in the Zou et al. (2026) simulations is calibrated to match low-redshift disk galaxies. This makes the comparison with the two observational samples consisting of galaxies at low redshift ($z\lesssim 0.1$) appropriate.
We thank the referee and Luis C. Ho for helpful comments and suggestions. Y.Z., F.Y., and S.J. are supported by the NSF of China (grants 12192220, 12192223, 12522301, 12133008, and 12361161601), the China Manned Space Program (grants CMS-CSST-2025-A08 and CMS-CSST-2025-A10), the National Key R&D Program of China No. 2023YFB3002502, and the National SKA Program of China (No. 2025SKA0130100). T. F. is supported by the National SKA Program of China No. 2025SKA0150103 and the National Natural Science Foundation of China under Nos. 12550002, 12133008, 12221003, and 11890692. We acknowledge the science research grants from the China Manned Space Project with grant Nos. CMS-CSST-2021-A04 and CMS-CSST-2025-A10. Numerical calculations were run on the CFFF platform of Fudan University, the supercomputing system at the Supercomputing Center of Wuhan University, and the High Performance Computing Resource at the Core Facility for Advanced Research Computing at Shanghai Astronomical Observatory.